# “How Big is Big Enough?” Adjusting Model Size in Continual Gaussian Processes

### v1

@@ Appendix D Proof of Guarantee
We cease the addition of points when $\widehat{\mathcal{U}}(M)-\widehat{\mathcal{L}}(M_{\mathbf{b}})<\alpha$. Since $\widehat{\mathcal{U}}(M)\geq\mathcal{L}^{*}$, then ^$-\widehat{\mathcal{L}}(M_{\mathbf{b}})<\alpha-\widehat{\mathcal{U}}(M)<\alpha-\mathcal{L}^{*}$. Assuming $\theta_{n}=\theta_{o}$, eq. (4) can be bounded as:
$\operatorname{KL}[q_{n}(f)\mid\mid p(f|\mathbf{y}_{o},\mathbf{y}_{n},\theta_{n})]$ $=\log p(\mathbf{y}_{n}|\mathbf{y}_{o})-\widehat{\mathcal{L}}+\Phi$ (24)
$\leq\log p(\mathbf{y}_{n}|\mathbf{y}_{o})+\alpha-\mathcal{L}^{*}+\Phi$
where $\Phi=-\int q_{n}(f)\log\frac{p(f|\mathbf{y}_{o},\theta_{o})}{q_{o}(f)}\text{d}f$. Let $q_{n}^{*}(f)=\mathcal{Z}^{-1}q_{o}(f)p(\mathbf{y}_{n}|f)$ be the variational distribution associated with =^$\mathcal{L}^{*}=\widehat{\mathcal{L}}(N_{n}+M_{\bf a})$, then by multiplying this value in both sides of the fraction inside the log, we obtain:
$\Phi$ $=\int q_{n}(f)\log\frac{q_{o}(f)}{p(f|\mathbf{y}_{o})}\text{d}f$ (25)
$=\int q_{n}(f)\log\frac{q_{o}(f)}{p(f|\mathbf{y}_{o})}\frac{q_{n}^{*}(f)}{q_{n}^{*}(f)}\text{d}f$
$=\int q_{n}(f)\log\frac{q_{o}(f)}{p(f|\mathbf{y}_{o})}\frac{q_{n}^{*}(f)}{\mathcal{Z}^{-1}q_{o}(f)p(\mathbf{y}_{n}|f)}\text{d}f$
$=\int q_{n}(f)\log\frac{q_{n}^{*}(f)}{p(f|\mathbf{y}_{n},\mathbf{y}_{o})p(\mathbf{y}_{n}|\mathbf{y}_{o})}\text{d}f+\log\mathcal{Z}$
$=\int q_{n}(f)\log\frac{q_{n}^{*}(f)}{p(f|\mathbf{y}_{n},\mathbf{y}_{o})}\text{d}f+\log\mathcal{Z}-\log p(\mathbf{y}_{n}\mid\mathbf{y}_{o})$
Then we can rewrite,
$\operatorname{KL}[q_{n}(f)\mid\mid p(f|\mathbf{y}_{o},\mathbf{y}_{n},\theta_{o})]$ (26)
$\quad\leq\log p(\mathbf{y}_{n}|\mathbf{y}_{o},\theta_{n})+\alpha-\mathcal{L}^{*}+\Phi$
$\quad\leq\log p(\mathbf{y}_{n}|\mathbf{y}_{o})+\alpha-\mathcal{L}^{*}+\int q_{n}(f)\log\frac{q_{n}^{*}(f)}{p(f|\mathbf{y}_{n},\mathbf{y}_{o},\theta_{n})}\text{d}f+\log\mathcal{Z}-\log p(\mathbf{y}_{n}\mid\mathbf{y}_{o})$
$\quad\leq\alpha+\int q_{n}(f)\log\frac{q_{n}^{*}(f)}{p(f|\mathbf{y}_{n},\mathbf{y}_{o},\theta_{n})}\text{d}f$
@@
$\operatorname{KL}[q_{n}(f)\mid\mid q_{n}^{*}(f)]$ $\leq\alpha$
Appendix E Inducing Points Selection Methods
The effectiveness of sparse approximations is determined by two key factors: the number of inducing points ($M$) and their locations. These parameters are responsible for a trade-off between the scalability and accuracy of the model, thus representing the captured information about the function $f$ [3, 24]. There exist multiple approaches to inducing point selection (see Burt et al. [3] for a review), broadly falling into two categories.
The first category includes choosing a subset of the input data through sampling or clustering and refining their location during gradient optimisation of the ELBO. For instance, Bui et al. [2] decided to sample a fixed number of inducing points from the set $\{Z_{o},X_{n}\}$ and optimise their locations. However, this approach uses a fixed memory, which requires a manual setting of the number of inducing points. Determining the memory size is inherently challenging as it will depend on the dataset’s properties and the estimated number of data points, not possible to know in advance.
In the second approach, points are iteratively selected from a set based on a specific criterion. Burt et al. [3] introduced an iterative "greedy variance" strategy that greedily chooses the location of the next inducing point to maximize the marginal variance in the conditional prior $p(f_{\neq\mathbf{u}}|\mathbf{u})$. This is equivalent to maximising $\operatorname{diag}[\mathbf{K_{ff}}-\mathbf{Q_{ff}}]$. The quantity $\operatorname{tr}\left(\mathbf{K_{ff}}-\mathbf{Q_{ff}}\right)$ is a bound for the norm of $\mathbf{K_{ff}}-\mathbf{Q_{ff}}$, and can therefore be used as a stopping criterion to determine the number of inducing points. We refer to the method that uses "greedy variance" with this stopping criterion as Conditional Variance. In this algorithm, new inducing points are no longer added once $\operatorname{tr}\left(\mathbf{K_{ff}}-\mathbf{Q_{ff}}\right)$ falls below a chosen tolerance value $\eta$. Although, an equivalent approach was studied in Fine and Scheinberg [12] and Foster et al. [13], this stopping criterion has not yet been tested in the literature. In the online setting, Chang et al. [4] use the “greedy variance” criterion by simply defining $\{\mathbf{Z}_{o},\mathbf{X}_{n}\}$ as the selection pool from which inducing point locations are selected, however, they maintain a fixed number of inducing points. Similarly, Maddox et al. [24] extends the “greedy variance” criterion to heteroskedastic Gaussian likelihoods and also uses a fixed memory approach. Galy-Fajou and Opper [14] introduced the Online Inducing Points Selection (OIPS) algorithm, a method that iteratively adds points from $\mathbf{X}_{n}$ to the existing set of inducing points: a point $\mathbf{x}\in\mathbf{X}_{n}$ is added to the set of inducing points if the maximum value of the covariance vector $\mathbf{k}_{\mathbf{u}}(\mathbf{x})$ is below a set threshold $\rho$. This method measures the potential potential impact of each new point with the existing inducing points set given the covariance threshold, this allows for an automatic selection of the number of inducing points unless a maximum is specified.
In our work, we are not interested in constraining the capacity of the model. Therefore, we will compare our method VIPS, to Conditional Variance with the selection pool ($\{\mathbf{Z}_{o},\mathbf{X}_{n}\}$) and OIPS without specifying a maximum number of inducing points. Table 2 shows a summary of the properties of these methods as well as the fixed memory approach used in Bui et al. [2].
Method Type Selection Pool Selection Criteria Stopping Criterion
Bui et al. [2] Fixed Random sample Gradient optimisation M constant
Cond. Variance (CV) Adaptive $\{X_{new},Z_{old}\}$ Greedy variance $\operatorname{tr}\left(\mathbf{K_{ff}}-\mathbf{Q_{ff}}\right)\leq\eta$.
OIPS [14] Adaptive $\{X_{new}\}$ $\max\mathbf{k}_{\mathbf{u}}(\mathbf{x})\leq\rho$ All points satisfying selection criteria.
VIPS Adaptive $\{X_{new}\}$ Greedy variance ^$\widehat{\mathcal{L}}\geq\widehat{\mathcal{U}}-\alpha$
Appendix F Further Experimental Details and Results
@@
F.1 Data types
@@
We use $N=500$ observations uniformly distributed from 00 to $10$. The data is ordered and divided into ten random batches.
@@
We simulate a scenario where small batches of data are received but the data is distributed across the input space. We use $N=150$ observations uniformly distributed from 00 to $10$. The data is shuffled and divided into ten random batches.
@@
We simulate a scenario where only outliers are encountered from time to time and the rest of the data is concentrated around a small part of the input space. We use two sets of data: the first set is sampled from a uniform distribution from $4$ to $6$, with $N=1000$ and the second set is sampled from a Cauchy distribution with a mean of $\mu=5$, with $N=300$. The data is divided into ten random batches, where the Cauchy observations are placed in the latter batches.
Figure 5: Plot of the three datasets considered in Section F.1.
F.2 Speed and Accuracy
F.2.1 Synthetic data
@@
For the synthetic dataset, we generate 1000 random noisy observations from the test function $f(x)=\sin(2x)+\cos(5x)$. We used a Squared Exponential kernel initialised with lengthscale $0.5$ and variance $1.5$. The noise variance was initialised to 0.5. The performance was measured on a test grid of $500$ points. For VIPS, $\delta=0.05$.
F.2.2 Timing
@@
This experiment was performed on an Nvidia RTX 6000’s GPU on a high-performance computing cluster. We used a Squared Exponential kernel with hyperparameters initialised to 1. The noise variance was initialised to 0.1. The dataset was divided into 20 batches, and we recorded the time in training per batch.
F.3 UCI datasets
Table 3: UCI datasets: methods parameters used for each dataset.
Dataset Dimension (N, D) Conditional Variance [3] OIPS [14] VIPS
$\eta$ $\rho$ $\delta$
Concrete 1030, 8 0.1 0.80 0.05
Skillcraft 3338, 19 0.1 0.80 0.05
Kin8nm 8192, 8 0.1 0.80 0.05
Elevators 16599, 18 0.1 0.80 0.05
Bike 17379, 17 0.1 0.80 0.05
These experiments were performed on an Nvidia RTX 6000’s GPU on a high-performance computing cluster. We used a Squared Exponential kernel with hyperparameters initialised to 1 for all datasets. The noise variance was initialised to 0.1. Table 3 shows the parameters used for each dataset and method. The smaller datasets $(<12000)$ were divided into 20 batches, and the larger $(>12000)$ into 50 batches. Figure 6 shows the results for all considered datasets. As mentioned in the main text, the method Conditional Variance often selected the most inducing points, leading to high accuracy. For most datasets, our methods had similar performance with fewer inducing points. Finally, OIPS selected the least inducing points, which led to a worse accuracy, especially in terms of the NLPD.
Figure 6: Plot of the mean number of inducing points, RMSE and NLPD over five 80%-train 20%-test random splits versus the number of data points, denoted by N, observed up to and including the current batch for different inducing point selection methods. The corresponding dataset is indicated at the left of the graph and each column corresponds to one of the metrics. The exact GP has access to all data up to and including the current batch.
Batch 1
Batch 5
Batch 10
Batch 15
Batch 20
Figure 7: A small robot with wheels is used to perform sequential estimation of magnetic field anomalies. Data is collected continuously as the robot moves along the path. Here the method Conditional Variance is used for the selection of inducing points.
Batch 1
Batch 5
Batch 10
Batch 15
Batch 20
Figure 8: A small robot with wheels is used to perform sequential estimation of magnetic field anomalies. Data is collected continuously as the robot moves along the path. Here the method VIPS is used for the selection of inducing points.

### v4

@@ Appendix C Proof of Guarantee
We cease the addition of points when $\widehat{\mathcal{U}}(M)-\widehat{\mathcal{L}}(M_{\mathbf{b}})<\alpha$. Since $\widehat{\mathcal{U}}(M)\geq\mathcal{L}^{*}$, then ^$-\widehat{\mathcal{L}}(M_{\mathbf{b}})<\alpha-\widehat{\mathcal{U}}(M)<\alpha-\mathcal{L}^{*}$. Eq.(4) can be bounded as:
$\operatorname{KL}[q_{n}(f)\mid\mid p(f|\mathbf{y}_{o},\mathbf{y}_{n},\theta_{n})]$ $=\log\frac{p(\mathbf{y}_{n},\mathbf{y}_{o}|\theta_{n})}{p(\mathbf{y}_{o}|\theta_{o})}-\widehat{\mathcal{L}}+\Phi$ (15)
$\leq\log\frac{p(\mathbf{y}_{n},\mathbf{y}_{o}|\theta_{n})}{p(\mathbf{y}_{o}|\theta_{o})}+\alpha-\mathcal{L}^{*}+\Phi$
where $\Phi=-\int q_{n}(f)\log\frac{p(f|\mathbf{y}_{o},\theta_{o})}{q_{o}(f)}\text{d}f$. Let $q_{n}^{*}(f)=\mathcal{Z}^{-1}q_{o}(f)p(\mathbf{y}_{n}|f)$ be the variational distribution associated with =^$\mathcal{L}^{*}=\widehat{\mathcal{L}}(N_{n}+M_{\bf a})$. Then, by expanding the true posterior and multiplying by the variational distributions $q_{n}^{*}(f)$ on both sides of the fraction inside the log, we obtain:
$\Phi$ $=\int q_{n}(f)\log\frac{q_{o}(f)}{p(f|\mathbf{y}_{o},\theta_{o})}\text{d}f$ (16)
$=\int q_{n}(f)\log\frac{q_{o}(f)p(\mathbf{y}_{o}|\theta_{o})}{p(\mathbf{y}_{o}|f)p(f|\theta_{o})}\text{d}f$
$=\int q_{n}(f)\log\frac{q_{o}(f)p(\mathbf{y}_{o}|\theta_{o})}{p(\mathbf{y}_{o}|f)p(f|\theta_{o})}\frac{q_{n}^{*}(f)}{q_{n}^{*}(f)}\text{d}f$
[…]
$=\int q_{n}(f)\log\frac{q_{n}^{*}(f)}{p(f|\mathbf{y}_{n},\mathbf{y}_{o},\theta_{o})}\text{d}f+\log\mathcal{Z}-\log\frac{p(\mathbf{y}_{n},\mathbf{y}_{o}|\theta_{o})}{p(\mathbf{y}_{o}|\theta_{o})}$
Using the above expansion for $\Phi$, eq. (15) becomes,
$\operatorname{KL}[q_{n}(f)\mid\mid p(f|\mathbf{y}_{o},\mathbf{y}_{n},\theta_{n})]$ (17)
$\quad\leq\log\frac{p(\mathbf{y}_{n},\mathbf{y}_{o}|\theta_{n})}{p(\mathbf{y}_{o}|\theta_{o})}+\alpha-\mathcal{L}^{*}+\Phi$
$\quad\leq\log\frac{p(\mathbf{y}_{n},\mathbf{y}_{o}|\theta_{n})}{p(\mathbf{y}_{o}|\theta_{o})}+\alpha-\mathcal{L}^{*}+\int q_{n}(f)\log\frac{q_{n}^{*}(f)}{p(f|\mathbf{y}_{n},\mathbf{y}_{o},\theta_{o})}\text{d}f+\log\mathcal{Z}-\log\frac{p(\mathbf{y}_{n},\mathbf{y}_{o}|\theta_{o})}{p(\mathbf{y}_{o}|\theta_{o})}.$
Assuming that $\theta_{n}=\theta_{o}$, the above can be simplified to
$\operatorname{KL}[q_{n}(f)\mid\mid p(f|\mathbf{y}_{o},\mathbf{y}_{n},\theta_{n})]\leq\alpha+\int q_{n}(f)\log\frac{q_{n}^{*}(f)}{p(f|\mathbf{y}_{n},\mathbf{y}_{o},\theta_{n})}\text{d}f$ (18)
@@
$\operatorname{KL}[q_{n}(f)\mid\mid q_{n}^{*}(f)]$ $\leq\alpha.$
Appendix D Further Experimental Details and Results
@@
D.1 Model size and data distribution
@@
We use $N=500$ observations uniformly distributed from 00 to $10$. The data is ordered and divided into ten batches.
@@
We simulate a scenario where small batches of data are received but the data is distributed across the input space. We use $N=150$ observations uniformly distributed from 00 to $10$. The data is shuffled and divided into ten batches.
@@
We simulate a scenario where only outliers are encountered from time to time and the rest of the data is concentrated around a small part of the input space. We use two sets of data: the first set is sampled from a uniform distribution from $4$ to $6$, with $N=1000$ and the second set is sampled from a Cauchy distribution with a mean of $\mu=5$, with $N=300$. The data is divided into ten batches, where the first few batches only contain observations from the $4$ to $6$ range, and the Cauchy observations are observed in the latter batches.
Figure 5: Plot of the three datasets considered in Section D.1.
D.2 The impact of model capacity in accuracy and training time
D.2.1 Accuracy comparison
@@
For the synthetic dataset, we generate 1000 random noisy observations from the test function $f(x)=\sin(2x)+\cos(5x)$. We used a Squared Exponential kernel initialised with lengthscale $0.5$ and variance $1.0$. The noise variance was initialised to $0.5$. The performance was measured on a test grid of $500$ points. For VIPS, we use $\delta=0.05$.
D.2.2 Training cost comparison
@@
This experiment was performed on an Nvidia RTX 6000’s GPU on a high-performance computing cluster. We used a Squared Exponential kernel with hyperparameters initialised to 1. The noise variance was initialised to 0.1. The dataset was divided into 20 batches, and we recorded the time in training per batch. For VIPS, we use $\delta=0.05$.
D.3 UCI datasets
These experiments were performed on an Nvidia RTX 6000’s GPU on a high-performance computing cluster. We used a Squared Exponential kernel with hyperparameters initialised to 1 for all datasets. The noise variance was initialised to 0.1. We consider six UCI (Dua & Graff, 2017) datasets of different characteristics: Concrete (1030, 8), Skillcraft (3338, 19), Kin8nm (8192, 8), Naval (11934, 14), Elevators (16599, 18), and Bike (17379, 17). The smaller datasets $(<12000)$ were divided into 20 batches, and the larger $(>12000)$ into 50 batches. The batches were created by sorting the data points by the first dimension.
Increasing the model size enhances performance until all relevant dataset information is captured; beyond this point, only computational costs increase. Therefore, when comparing inducing point methods, it is essential to consider the size-performance trade-off, rather than focusing solely on performance gains. This trade-off is typically controlled by a hyperparameter of the model. However, in continual learning, traditional cross-validation for tuning is not practical since it would require storing all past data. As a result, a good method needs to use pre-set hyperparameters and still perform well on different datasets. We compare our method, VIPS, with two other adaptive approaches: Conditional Variance (CV) and OIPS. Our goal is to determine if a single hyperparameter for each method can perform consistently well across different datasets. To do this, we evaluate various hyperparameter settings for each method, with ranges $\delta\in[0.005,0.2]$ for VIPS, $\delta\in[0.005,1.0]$ for CV, $\rho\in[0.80,0.999]$ for OIPS.
For each method, we identified hyperparameter values that achieved a particular accuracy threshold (either RMSE or NLPD) within a percentage of 5% or 10% of the full-batch GP relative to a noise model. For example, for 5% RMSE, the accuracy threshold is calculated as: $\mathrm{RMSE}_{\mathrm{exact}}+0.05\left|\mathrm{RMSE}_{\mathrm{noise}}-\mathrm{RMSE}_{\mathrm{exact}}\right|$. Among the hyperparameter values meeting these criteria, we select the one that results in the smallest model size. For CV and VIPS, this corresponds to the largest hyperparameter value, while for OIPS, it corresponds to the smallest hyperparameter value (see Table 5 for a summary of the methods).
With the optimal hyperparameter selected, all methods achieve the desired performance threshold. Therefore, we can compare the methods based on their model size, where smaller sizes are preferred. As explained in the main part of the paper, we did not find a hyperparameter that satisfied the accuracy thresholds for OIPS in the Naval dataset. For this method, we selected the hyperparameter configuration that satisfied our conditions for five out of six datasets, treating Naval as an outlier. This decision was made considering that OIPS already selected larger models with this less stringent hyperparameter (for example 80% of the data points for the Kin8nm dataset) and extending the hyperparameter range would have led to even larger dataset sizes.
Since such dataset-specific limitations are common in practical applications, and to ensure a fair comparison across methods, we evaluated configurations that allowed for either one or two datasets to be outliers for each method. To report performance, we calculated the difference with the full-batch GP metric relative to a noise model. For each dataset and batch, we computed the relative RMSE percentage as
$\mathrm{RMSE}\,\%=\left(\frac{\mathrm{RMSE}_{\mathrm{method}}-\mathrm{RMSE}_{\mathrm{exact}}}{\left|\mathrm{RMSE}_{\mathrm{noise}}-\mathrm{RMSE}_{\mathrm{exact}}\right|}\right)\times 100$
and similarly for the NLPD metric. The test set for each batch consists only of data from the current and previous batches.
Figure 6 shows a summary of the results, presenting the absolute counts for smallest model size. Tables 6, 7, 8 and 9 provide details for each threshold and outlier choice. The results reveal consistent patterns across different accuracy thresholds and metrics. VIPS achieves the smallest model size while maintaining performance within thresholds, particulary when all datasets are considered.
Allowing outlier datasets reduces model sizes across all methods. This comes with performance degradation on the particular outlier. We observe that this degradation is often higher for CV and OIPS. The Naval dataset proved particularly challenging for these methods. This is probably due to its almost noiseless nature. While VIPS came close (missing the threshold by 0.48%), none of the methods achieved the 5% target NLPD threshold within the tested hyperparameter range. Given the dataset’s near-zero noise characteristics, we chose not to expand the hyperparameter range to avoid excessive model sizes.
The results for OIPS and CV suggest that their hyperparameters may need to be tuned for each dataset. As mentioned, this is not possible in continual learning where data is not available in advance. In contrast, across all configurations, VIPS demonstrates consistent behaviour, minimising model size in the majority of cases and hence reducing the need for dataset-specific tuning. Based on our experimental results, we recommend using $\delta=0.035$ for VIPS, which represents the average of successful parameter values when all datasets were considered. This value should provide near-exact performance across different scenarios while minimising model size.
Figure 6: Number of datasets where each method achieves minimal model size at different outlier allowances. A “win” is assigned to the dataset with the smallest model which satisfies the accuracy thresholds. Counts represent absolute wins. Higher counts indicate better method robustness across datasets.
Table 6: Mean (std) of inducing points and RMSE % difference with full-batch GP for the final batch at different outlier allowances at the selected operating point to achieve a 5% RMSE threshold. ✗ Gray highlights cases that exceed the threshold, and bold text identifies the smallest model size. ‘Max.’ denotes reaching the model’s maximum capacity limit. For none, one, and two outliers, the selected hyperparameters were: VIPS $\delta=0.02,0.065,0.14$, CV $\eta=0.005,0.06,0.5$ and OIPS $\rho=-,0.955,0.935$.
Dataset Outliers VIPS CV OIPS
M RMSE % M RMSE % M RMSE %
Concrete None 434(79) 3.67% 492(54) 3.28% - -
One 317(101) 4.24% 321(71) 4.65% 383(90) 4.89%
Two 163(95) 10.43% ✗ 70(5) ✗  13.34% ✗ 257(106) ✗ 8.84%
Skillcraft None 177(13) 0.14% 739(34) 0.02% - -
One 135(1) 0.21% 312(15) 0.04% 236(38) 0.07%
Two 133(0) 0.13% 130(56) 1.45% 185(21) 0.02%
Kin8nm None 4281(69) 0.15% 6316(9) 0.11% - -
One 2513(54) 1.35% 5685(26) 0.13% 6523(9) 0.12%
Two 1356(42) 4.97% 5334(27) 0.13% 6499(10) 0.12%
Naval None 373(19) 0.40% 49(1) 2.39% - -
One 86(3) 1.37% ✗ 23(7) ✗ 79.08% ✗  20(1) ✗ 407.22%
Two 43(2) 2.67% ✗ 21(5) ✗ 336.08% ✗ 20(1) ✗ 407.22%
Elevators None 438(17) 1.19% 3133(140) 0.31% - -
One 312(5) 1.33% 1194(42) 0.85% 386(47) 1.19%
Two 276(4) 1.55% 536(22) 1.43% 305(13) 1.49%
Bike None 2296(138) 4.61% Max. 7000 0.58% - -
One ✗ 853(19) ✗  8.73% 5563(36) 1.02% 3916(122) 2.19%
Two ✗ 498(10) ✗ 11.24% 3940(41) 1.80% 2438(130) 4.23%
Table 7: Mean (std) of inducing points and RMSE % difference with full-batch GP for the final batch at different outlier allowances at the selected operating point to achieve a 10% RMSE threshold. ✗ Gray highlights cases that exceed the threshold, and bold text identifies the smallest model size. ‘Max.’ denotes reaching the model’s maximum capacity limit. For none, one, and two outliers, the selected hyperparameters were: VIPS $\delta=0.095,0.12,0.2$, CV $\eta=0.005,0.13,0.6$ and OIPS $\rho=-,0.93,0.845$.
Dataset Outliers VIPS CV OIPS
M RMSE % M RMSE % M RMSE %
Concrete None 234(116) 7.66% 492(54) 3.28% - -
One 178(95) 9.81% 184(74) 9.96% 240(102) 9.03%
Two ✗ 111(37) ✗ 12.93% ✗ 3(2) ✗ 126.81% ✗ 108(66) ✗ 20.79%
Skillcraft None 134(0) 0.12% 739(34) -0.02% - -
One 133(0) 0.13% 226(22) 0.20% 174(16) 0.04%
Two 133(0) 0.12% 17(0) 7.93% 138(1) 0.06%
Kin8nm None 1904(51) 2.55% 6316(9) 0.11% - -
One 1563(39) 3.77% 5334(27) 0.13% 6458(8) 0.11%
Two 949(25) 8.17% 3408(32) 0.19% 5053(93) 0.14%
Naval None 57(1) 2.25% 49(1) 2.39% - -
One 48(3) 2.57% ✗ 23(8) ✗ 323.90% ✗ 16(1) ✗ 315.28%
Two 38(1) 5.07% ✗ 19(3) ✗ 553.11% ✗ 13(1) ✗ 407.6%
Elevators None 291(4) 1.41% 3133(140) 0.31% - -
One 281(4) 1.51% 836(27) 1.11% 298(8) 1.49%
Two 267(2) 1.57% 287(18) 2.52% 269(1) 1.59%
Bike None 650(17) 9.98% Max. 7000 0.58% - -
One ✗ 549(14) ✗ 10.82% 4895(28) 1.29% 1964(137) 5.18%
Two ✗ 400(7) ✗ 12.50% 2582(66) 4.22% 594(11) 10.00%
Table 8: Mean (std) of inducing points and NLPD % difference with full-batch GP for the final batch at different outlier allowances at the selected operating point to achieve a 5% NLPD threshold. ✗ Gray highlights cases that exceed the threshold, and bold text identifies the smallest model size. ‘Max.’ denotes reaching the model’s maximum capacity limit. For one, and two outliers, the selected hyperparameters were: VIPS $\delta=0.005,0.015$, CV $\eta=0.025,0.100$ and OIPS $\rho=0.975,0.960$.
Dataset Outliers VIPS CV OIPS
M NLPD % M NLPD % M NLPD %
Concrete One 514(74) 2.27% 417(63) 4.28% 490(73) 4.72%
Two 451(77) 4.43% ✗ 227(84) ✗ 15.84% ✗ 410(88) ✗ 8.34%
Skillcraft One 291(34) 0.60% 439(29) 0.60% 488(119) 0.53%
Two 195(18) 0.78% 260(16) 0.81% 332(82) 0.51%
Kin8nm One 5640(73) 0.85% 6194(12) 0.88% 6547(8) 0.88%
Two 5065(46) 0.94% 5334(27) 0.90% 6539(8) 0.88%
Naval One ✗ 1389(18) ✗ 5.48% ✗ 27(5) ✗ 4654.12% ✗ 26(1) ✗ 115.29%
Two ✗ 509(15) ✗ 9.99% ✗ 28(4) ✗ 8146.44% ✗ 20(1) ✗ 133.87%
Elevators One 740(24) 1.74% 1814(88) 1.23% 1869(257) 1.34%
Two 487(24) 2.02% 916(21) 1.91% 642(133) 2.01%
Bike One 5064(334) 3.99% Max. 7000 2.13% 6794(79) 2.55%
Two ✗ 2794(148) ✗ 8.18% 4895(28) 3.54% 5131(64) 4.05%
Table 9: Mean (std) of inducing points and NLPD % difference with full-batch GP for the final batch at different outlier allowances at the selected operating point to achieve a 10% NLPD threshold. ✗ Gray highlights cases that exceed the threshold, and bold text identifies the smallest model size. ‘Max.’ denotes reaching the model’s maximum capacity limit. For one, and two outliers, the selected hyperparameters were: VIPS $\delta=0.015,0.02,0.05$, CV $\eta=0.001,0.07,0.5$ and OIPS $\rho=-,0.95,0.935$.
Dataset Outliers VIPS CV OIPS
M NLPD % M NLPD % M NLPD %
Concrete None 451(77) 4.43% 492(54) 2.32% - -
One 434(79) 5.68% 303(78) 9.64% 383(90) 9.56%
Two 349(86) 9.04% ✗ 3(2) ✗ 118.71% ✗ 247(101) ✗ 15.73%
Skillcraft None 195(18) 0.78% 1091(107) -0.34% - -
One 177(13) 0.78% 295(13) 0.71% 236(38) 0.57%
Two 141(4) 1.10% ✗ 54(45) ✗ 8.03% 185(21) 0.68%
Kin8nm None 5065(46) 0.94% 6469(7) 0.87% - -
One 4281(69) 0.34% 5685(26) 0.91% 6523(9) 0.88%
Two 2953(71) 1.89% 5334(27) 0.90% 6499(10) 0.88%
Naval None 509(15) 9.99% 903(8) 6.17% - -
One ✗ 373(19) ✗ 11.92% ✗ 27(6) ✗ 2405.87% ✗ 20(1) ✗ 133.87%
Two ✗ 126(6) ✗ 22.95% ✗ 22(7) ✗ 6012.82% ✗ 20(1) ✗ 133.87%
Elevators None 487(24) 2.02% 3133(140) 0.82% - -
One 438(17) 2.11% 1099(50) 1.73% 386(47) 2.68%
Two 332(8) 2.54% 352(29) 3.76% 305(13) 4.84%
Bike None 2794(148) 8.18% Max. 7000 2.04% - -
One 2296(138) 9.92% 5563(36) 2.95% 3916(122) 5.50%
Two ✗ 1037(23) ✗ 17.16% 2582(66) 9.61% 2438(130) 9.66%
D.4 Magnetic anomalies
The data used in this experiment is obtained from Solin et al. (2018) and is available on GitHub. The objective of this task is to detect local anomalies in the Earth’s magnetic field online, caused by the presence of bedrock and magnetic materials in indoor building structures. For this purpose, a small robot with a 3-axis magnetometer moves around an indoor space of approximately 6 meters by 6 meters and measures the magnetic field strength. Out of the 9 available trajectories, we use trajectories 1, 2, 3, 4, and 5 (with $n=8875,9105,9404,7332,8313$, respectively) for the experiments. Specifically, we use trajectories 1, 2, 4 and 5 for the first experiment and trajectory 3 for the second.
We use the experimental setup proposed in Chang et al. (2023a). The proposed model applies a GP prior to magnetic field strength, given by $\mathcal{G}\mathcal{P}\left(0,\sigma_{0}^{2}+\kappa_{\sigma^{2},\ell}^{\mathrm{Mat}}\left(\mathbf{x},\mathbf{x}^{\prime}\right)\right)$ (in $\mu\mathrm{T}$ ), where the kernel consist of a constant kernel and a Matérn-$\nu/2$ kernel. The model assumes the spatial domain is affected by Gaussian noise with a variance $\sigma_{\mathrm{n}}^{2}$. The initial variance for the constant kernel is set to $500$, and the Gaussian likelihood is initialised with a noise variance of $0.1$.
Our aim for this experiment is to test the optimal hyperparameters identified in the previous section for the 10% RMSE threshold for each adaptive method in a real-world setting. The setting simulates an ever-expanding domain, where the robot is not confined to a predefined area. In this context, the model continuously learns new parts of the space. Therefore, a method that works will need to sufficiently expand the model’s size to accommodate new data without letting it grow uncontrollably.
In the first experiment, we aim to sequentially learn the paths taken by the robot using trajectories 1, 2, 4 and 5, i.e. an entire path will correspond to one batch. We investigate whether the method can adapt to changes in the environment and adjust the number of inducing points accordingly. During this process, we concurrently learn the hyperparameters $\sigma_{0}^{2}$, $\sigma^{2}$, $\ell$, and $\sigma_{\mathrm{n}}^{2}$. As a test set, we use trajectory 3. Figures 7, 9, 11 show the temporally updating field estimate over batches alongside the corresponding path travelled in each batch.
In the second experiment, we focus on the streaming learning of trajectory 3. The trajectory is split into 20 batches. We compare the number of inducing points selected and the estimate obtained by the three methods, Conditional Variance (CV), OIPS and VIPS (ours). Detailed learning of the path for each method is shown in figures 8, 10 and 12. As a test set, we use trajectories 1, 2, 4, and 5. We observed how Conditional Variance chooses an excessive number of inducing points, indicating that its hyperparameter needs tuning, which is impractical in the continual learning setting. OIPS chooses the least inducing points; as seen in the figure, the number of inducing points is more concentrated at the beginning of the path and becomes sparse towards the end. When compared to the learning estimates derived from the paths in the previous experiment, we see that both estimates differ significantly, suggesting that OIPS fails to adequately add enough capacity to capture the changes in the environment. VIPS provides the middle ground, selecting a moderate number of inducing points that effectively balance accuracy and memory size. This choice allows VIPS to maintain a robust estimate of the magnetic field obtained when compared to learning by paths without excessive computational overhead.
In the last two sections, when compared to both alternative approaches, VIPS achieved the best trade-off between performance and model size without requiring hyperparameter tuning, making it the preferred method among the three.